# Randomly Sampled Language Reasoning Problems Reveal Limits of LLMs

### v2

@@ A.1 Sampling of DFAs
We use rejection sampling to sample DFAs. Specifically, we uniformly sample a start state, then for each (source state, symbol) pair, we sample a post-transition state. We also randomly assign each state to be accept or reject with probability 50%. We then reject any DFA that has all accept or all reject states, or for which certain states are unreachable from the start state.
@@ A.2 Sampling of Sequence Completion Tasks
To sample a task instance, we sample example sequences and distinct prefix. Each example sequence is sampled uniformly from the space of $\{a,b,c\}^{10}$ and then rejected if the DFA does not accept the sequence. Our distinct prefix and completion are sampled uniformly from $\{a,b,c\}^{5}\times\{a,b,c\}^{5}$, and is rejected if the DFA does not accept the concatenation of the two, or if the prefix is the prefix of any of the previous sequences. We then discard the completion. If we, at any point, reject 50 sequences when attempting to sample a sequence or prefix, we return an error.
We run a “pilot” sampling for a DFA to ensure that it is valid, in which we sample an instance as described abov. If there is an error in sampling this pilot instance, we reject the DFA. Otherwise, we proceed to sample our task instances. At this stage, if there is an error in sampling, we reject the instance rather than the DFA.
@@ A.3 Sampling of Transducer Tasks
Appendix B Sequence Completion task prompt with Commas
@@
Appendix C More details on Sum Modulo 3 DFA case study
Figure 4: Results on Sum Modulo 3 DFA under trivial / nontrivial conditions. Percentages are accuracy numbers, and venn diagram is error counts. (a) In this condition, CR and the 6-GramT both get very high accuracies, with nearly all 6-GramT also being CR errors. QB does poorly. (b) In this condition, models do significantly more poorly overall, with CR in particular performing worse than chance. Here, errors are more symmetric, with more 6-GramT errors that are not accounted for by either or both model, indicating that a larger fraction of both successes and failures in this condition are down to random chance.
Appendix D Prompt Listings
@@ D.1 Summaries
D.2 Full example listings
D.2.1 BasicT

### v3

@@ A.1 Sampling of DFAs
We use rejection sampling to sample DFAs. Specifically, we uniformly sample a start state, then for each (source state, symbol) pair, we sample a post-transition state. We also randomly assign each state to be accept or reject with probability 50%. We then reject any DFA that has all accept or all reject states (so only DFAs with 1 or 2 accept states are allowed), or for which certain states are unreachable from the start state.
@@ A.2 Sampling of Sequence Completion Tasks
To sample a task instance, we sample example sequences and distinct prefix. Each example sequence is sampled uniformly from the space of $\{a,b,c\}^{10}$ and then rejected if the DFA does not accept the sequence. Our distinct prefix and completion are sampled uniformly from $\{a,b,c\}^{5}\times\{a,b,c\}^{5}$, and are rejected if the DFA does not accept the concatenation of the two, or if the prefix is the prefix of any of the previous sequences. We then discard the completion. If we, at any point, reject 50 sequences when attempting to sample a sequence or prefix, we return an error.
We run a “pilot” sampling for a DFA to ensure that it is valid, in which we sample an instance as described above. If there is an error in sampling this pilot instance, we reject the DFA. Otherwise, we proceed to sample our task instances. At this stage, if there is an error in sampling, we reject the instance rather than the DFA. This pilot sample rejection procedure leads to a slight bias towards 2-accept state DFAs over 1-accept state DFAs, as measured by the $\textsc{Random}_{S}$ baseline.
@@ A.3 Sampling of Transducer Tasks
Appendix B Results of o1-preview
We evaluated o1-preview on 10 DFAs, using 30 problem instances per DFA of the Transducer task, as in other Transducer experiments, and the $\textsc{Basic}_{T}$ prompt, as this is the most neutral prompt. Table 3 displays results on each DFA. Overall, while these results are not on a particularly large sample, they fairly definitively demonstrate that o1-preview does not achieve strikingly good performance on this task.
DFA o1-preview gpt-4o 6-Gram
1 25/30 27/30 26/30
2 23/29 24/30 25/30
3 19/30 23/30 28/30
4 22/30 23/30 28/30
5 29/29 30/30 30/30
6 19/30 24/30 30/30
7 17/29 23/30 25/30
8 23/30 25/30 26/30
9 21/30 28/30 30/30
10 29/30 29/30 30/30
Table 3: Results on each DFA. We find that in 7 cases, o1-preview underperforms gpt-4o, in 2 cases it gets the same number of instances wrong but provides a non-answer on an additional instance, and in 1 case it ties gpt-4o. In no cases does it outperform.
Appendix C More details on Sum Modulo 3 DFA case study
Figure 5 depicts the results of the Sum Modulo 3 experiment, but filtered for two conditions. In the (a) condition, the trace ends in such a way that a 3-Gram model would be able to determine the output, and the (b) condition is the complement.
Figure 5: Results on Sum Modulo 3 DFA under trivial / nontrivial conditions. Percentages are accuracy numbers, and venn diagram is error counts. (a) In this condition, CR and the 6-GramT both get very high accuracies, with nearly all 6-GramT also being CR errors. MB does relatively poorly. (b) In this condition, models do significantly more poorly overall, with CR in particular performing worse than chance. Here, errors are more symmetric, with more 6-GramT errors that are not accounted for by either or both model, indicating that a larger fraction of both successes and failures in this condition are down to random chance.
@@ Appendix C More details on Sum Modulo 3 DFA case study
Appendix D Sequence Completion task prompt with Commas
@@
Appendix E Prompt Listings
@@ E.1 Summaries
E.2 Full example listings
E.2.1 BasicT